The second angle in a triangle is 5 times as large as the first, the third angle is 2/3 as large than the first
Answer by www.mathmartian.com(5)

The angles in a triangle add up to 180 degrees.
Let's name the first angle x.
The second angle is 5 times as large as the first, or 5x.
The third angle is 2/3 as large as the first, or 2/3x.
The sum of all three angles should be 180, so let's write an equation and solve for x.
x+5x+2/3x=180
6 2/3x=180
20/3x=180 convert 6 2/3 to an improper fraction
*3/20 *3/20
x=27
So the first angle is 27 degrees
5x=5*27=135
The second angle is 135 degrees
2/3x=2/3*27=18
The third angle is 18 degrees
Let's check our answer. All three angles should add up to 180 degrees.
27+135+18=180
Yes, our answer checks out:)